# PersoBench: Benchmarking Personalized Response Generation in Large Language Models

### v1

@@ 4.2 Implementation setup
Post-processing Instructing LLMs to generate output in JSON format (Table 1), any unparsable outputs are considered failed instances and factored into the failure ratio calculation. From the successfully parsed outputs, we extracted only the generated response to compare against the golden response, discarding the reasoning portion, as the analysis focused on the impact of COT reasoning on the response itself. Additionally, response generation time, measured in seconds, provides valuable insights for practical industrial applications, particularly when deploying open-source LLMs.
Metrics: We evaluate the generated responses across fluency, diversity, coherence, and personalization. Fluency is measured using ROUGE [26], METEOR [27], and BERTScore [28] to assess linguistic quality and semantic similarity. Diversity is evaluated using Dist-1 and Dist-2 [29] to calculate the ratio of unique unigrams and bigrams. For personalization, the Consistency Score (C Score) [30] measures persona consistency, while the Persona Distance (P-Dist) score [31] assesses persona coverage. Coherence is evaluated with the Utterance Entailment Score (UE-Score) [32] based on Natural Language Inference (NLI). These metrics ensure a comprehensive evaluation of the model’s performance.
@@
LLMS: To implement an automatic evaluation pipeline, we chose the instructed versions of LLMs. This allows us to automate content extraction from the generated responses, streamlining the post-processing and evaluation modules. In this regard, we selected Mistral 7B [^4] https://mistral.ai/news/announcing-mistral-7b/, Qwen2 7B [^5] https://github.com/QwenLM/Qwen2, and Llamma3.1 8B [^6] https://llama.meta.com/ as the three open-source LLMs. For the closed-source models, we utilized the GPT series from OpenAI [^7] https://openai.com/research/gpt, including GPT-3.5 Turbo, GPT-4 Turbo, and GPT-4o Mini.
4.3 Expriment Results
Table 5: Fluency evaluation of LLMs on the FoCus dataset.
Model Vanilla Setup COT Setup
ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1) ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1)
Mistral 7B 0.11/0.03/0.10 0.11 0.48/0.49/0.48 0.15/0.03/0.13 0.15 0.75/0.77/0.76
Qwen2 7B 0.08/0.02/0.08 0.10 0.41/0.43/0.42 0.10/0.02/0.09 0.13 0.60/0.63/0.62
Llama3.1 8B 0.11/0.03/0.10 0.11 0.47/0.47/0.47 0.15/0.03/0.13 0.16 0.80/0.81/0.80
GPT3.5 Turbo 0.20/0.06/0.18 0.20 0.78/0.79/0.78 0.17/0.04/0.15 0.18 0.79/0.80/0.80
GPT4o Mini 0.19/0.06/0.17 0.22 0.78/0.81/0.80 0.20/0.05/0.17 0.22 0.85/0.87/0.86
GPT4 Turbo 0.23/0.08/0.20 0.23 0.84/0.84/0.84 0.22/0.07/0.19 0.23 0.85/0.86/0.86
@@ 4.3 Expriment Results
Table 6: Diversity evaluation of LLMs using Dist-1 and Dist-2 metrics across benchmarking datasets in Vanilla and COT setups.
Model IT-ConvAI2 Blended Skill Talk FoCus
@@
Can LLMs generate fluent responses? The benchmarked LLMs generally demonstrate strong fluency across various datasets, regardless of context length. This is particularly evident in Tables 3, 4 and 5, in the high BERTScore values, which show that the models effectively capture the meaning and structure of conversations. While ROUGE and METEOR scores indicate less reliance on exact phrase matches with the golden responses, this does not detract from overall fluency. BERTScore proves to be a more practical metric in conversational AI, where capturing the intended meaning is more important than exact word matches.
Both groups perform well when comparing open-source and closed-source LLMs, with closed-source models like those in the GPT-4 series slightly outperforming their open-source counterparts, particularly in more extended contexts such as those in the FoCus dataset. However, the performance gap is not substantial, showing that open-source models are still capable, especially in less complex dialogues.
@@
Among the open-source LLMs, Qwen2 performs well in shorter contexts, while Llama3.1 excels in longer and more challenging dialogues, such as the FoCus dataset. Llama3.1’s larger parameter size and effective use of COT reasoning allow it to handle more complex contexts, surpassing other open-source models. Meanwhile, GPT-4 models consistently demonstrate robust fluency across all contexts. Although COT generally enhances their fluency in shorter contexts, it can sometimes introduce unnecessary complexity, slightly reducing performance by making responses more rigid and formulaic. Despite this, GPT-4 models maintain strong overall fluency in both short and long dialogues.
@@
How proficient are LLMs at generating diverse responses? Benchmarked LLMs are generally good at generating diverse responses, particularly in shorter-context datasets like IT-ConvAI2 and Blended Skill Talk (Table 6). This is supported by high Dist-1 and Dist-2 scores, which measure the diversity of unigrams and bigrams, indicating a wide variety of words and varied word combinations in the generated responses. When comparing the closed-source and open-source LLMs, closed-source models like those in the GPT-4 series tend to perform slightly better overall, especially in longer contexts. However, the difference between the two groups is not stark, indicating that open-source models are still quite capable, particularly in less complex dialogues.
For open-source LLMs, performance tends to decline as the context length increases, as seen in the FoCus dataset, which involves longer and more complex dialogues. This suggests that while these models handle diversity well in shorter exchanges, they face challenges maintaining the same diversity level in longer contexts.
The impact of COT is particularly noticeable in open-source LLMs when dealing with longer contexts. For instance, Llama3.1 benefits the most from COT, likely due to its larger parameter size, which helps manage more intricate dialogues, as it has an extra billion parameters compared to other open-source LLMs. On the other hand, closed-source LLMs like those from the GPT-4 series show minimal improvement with COT. In some cases, especially in shorter contexts, their performance even slightly degrades. This may occur because COT can cause these models to overstructure their responses, leading to less natural diversity and producing too rigid or formulaic outputs. Despite this, the GPT-4 generation consistently stands out as top performers across varying context lengths, demonstrating robust diversity generation in both short and long dialogues.
@@
Table 7: Coherence evaluation using UE Score on benchmarking datasets in Vanilla and COT setups.
Model IT-ConvAI2 Blended Skill Talk FoCus
@@
Mistral 7B 0.46 0.31 0.48 0.44 0.21 0.33
Qwen2 7B 0.42 0.38 0.39 0.52 0.17 0.36
Llama3.1 8B 0.23 0.22 0.33 0.37 0.17 0.33
GPT 3.5 Turbo 0.49 0.38 0.51 0.57 0.37 0.45
GPT 4o Mini 0.40 0.38 0.38 0.39 0.35 0.37
GPT4 Turbo 0.43 0.42 0.56 0.50 0.39 0.45
Table 8: Personalization evaluation of LLMs using P-Dist Score and C Score across the benchmarking dataset, in Vanilla and COT setups.
Model IT-ConvAI2 Blended Skill Talk FoCus
Vanilla COT Vanilla COT Vanilla COT
P-Dist Score C Score P-Dist Score C Score P-Dist Score C Score P-Dist Score C Score P-Dist Score C Score P-Dist Score C Score
@@
Can LLMs generate responses that maintain coherence with the given context and persona? Based on the UE score presented in Table 7, which is calculated using an NLI model whic relies on natural language inference, LLMs generally struggle with generating coherent responses, especially when working with long-context datasets like FoCus. Even in shorter-context datasets, as seen in the IT-ConvAI2 results, LLMs show poor performance in maintaining coherence. While the closed-source LLMs perform better at this task, even powerful models like GPT-4-Turbo fall short of achieving good results. Among them, GPT4o Mini demonstrates marginally lower performance across all the datasets with varying context sizes. While COT reasoning offers a modest performance improvement, GPT series LLMs, as fine-tuned LLMs, remain limited in their ability to handle the coherence task effectively.
Although the open-source LLMs’ performances are closely aligned, Llama3.1 exhibits marginally lower results across the datasets, indicating its inability to leverage the additional billion parameters effectively. It also benefits less from COT compared to the others. Overall, open-source LLMs face significant challenges in maintaining coherence, particularly in complex, long-context scenarios.
@@
Can LLMs generate personalized responses? The results of P-Dist and C scores, which are shown in Table 8, two well-established NLI-based metrics, reveal that LLMs struggle significantly with personalization. There is no clear benefit in using closed-source LLMs over open-source alternatives, underscoring the inherent difficulty of the personalization task for these models. Even in the best scenarios, the P-Dist score—where 1 represents the ideal outcome—demonstrates only moderate performance. The C score, which also penalizes contradictory entailments with negative values, exposes these limitations even more starkly, particularly in short-length contexts where less content is available for personalization. On short-context datasets, no clear performance trend emerges among open-source LLMs. However, on the FoCus dataset, Llama3.1 shows a slight superiority in the personalization task. The same overall scenario applies for close-source LLMs, while GPT-4 Turbo emerges as the least effective performer. COT negatively impacts short-length conversations but yields a notable improvement in long-length conversations, as seen with the FoCus dataset. This suggests that the added complexity of COT is only advantageous when there is enough content to work with. This need for COT reasoning aligns with findings from [33], which highlight that current language models do not reliably utilize information in long input contexts.
5 Analysis and Discussion
To further analyse the LLM performance on longer text, we focused on the FoCus dataset. Our previous results show that its long-context nature makes LLM performance differences more apparent. Fig. 2 provides an overall view of LLM groups across different prompting configurations. BertScore-F1, Dist-2, UE Score, P-Dist, and C Score are considered for overall fluency, diversity, coherence, persona coverage and persona consistency analysis. Response personalization remains challenging, even for fine-tuned closed-source LLMs, as evidenced by the results in both persona consistency and persona coverage. Additionally, the coherence evaluation, which accounts for both persona and golden response alignment, further highlights this difficulty. While open-source LLMs generally cannot compete with closed-source models in terms of fluency and diversity, their performance in personalization and coherence is considerably enhanced with the use of COT. In fact, an open-source model like Llama3.1 can even surpass fine-tuned closed-source LLMs in these areas. Meanwhile, the performance of closed-source LLMs do not show any tangible improvement.
@@ 5 Analysis and Discussion
Figure 2: Performance analysis of (a) open-source LLMs in vanilla setting (b) close-source LLMs in vanilla setting (c) open-source LLMs in COT setting (b) close-source LLMs in COT setting on the FoCus dataset.
@@
Figure 3: Analysis of response time and failure ratio for LLMs on the FoCus dataset: (a) Vanilla setup and (b) COT setup.
Fig.3 illustrates LLMs’ comparative performance from response time and failure ratio, two aspects that can provide information for practical insights for selecting LLMs for industrial use. The failure ratio can indicate the model’s instructability, which refers to how well the model follows instructions, particularly considering the maximum token limit provided during response generation. While COT increases response time, it directly enhances the instructability of both open-source and closed-source LLMs, with a significant difference in their ability to handle varying token length limits.

### v2

@@ 4.1 Implementation setup
As we develop our automatic benchmarking framework necessitates instructing LLMs to generate output in JSON format (Table 1), any unparsable outputs are considered failed instances and factored into the failure ratio calculation. However, the raw generated responses are stored and made available in the benchmark’s GitHub repository, allowing researchers to conduct further analysis. From the successfully parsed outputs, we extracted only the generated response to compare against the golden response, discarding the reasoning portion, as the analysis focused on the impact of CoT reasoning on the response itself. Additionally, response generation time, measured in seconds, provides valuable insights for practical industrial applications, particularly when deploying open-source LLMs.
Metrics: We evaluate the generated responses across fluency, diversity, coherence, and personalization. Metrics like fluency and diversity assess fundamental aspects such as grammatical accuracy, relevance, and clarity. These qualities are essential for any dialogue system, as personalization holds no value if the responses themselves are not high-quality. Fluency is measured using BERTScore [43] and the Naturalness dimension of UniEval framework [45] to assess linguistic quality and semantic similarity. Diversity is evaluated using Dist-1 and Dist-2 [24] to calculate the ratio of unique unigrams and bigrams. Building on these foundational qualities, our automated benchmarking pipeline specifically evaluates personalization and coherence as its main focus. These aspects assess a system’s ability to adapt responses to user preferences while maintaining contextual consistency. We selected these criteria based on a recently published survey [10], reflecting current research priorities in dialogue systems. While these metrics have been explored individually in prior studies, our work is the first to systematically benchmark LLMs with a dedicated focus on these specific aspects.
For personalization, we evaluate model behavior from two complementary perspectives: persona consistency and persona coverage, each measured using a tailored metric. Persona consistency reflects whether the generated response logically aligns with the given persona traits. To measure this, we use the Consistency Score (C-score) [32], which applies a Natural Language Inference (NLI) model fine-tuned on the Dialogue NLI (DNLI) dataset. It classifies the relationship between each persona sentence and the response as entailment, contradiction, or neutral, and aggregates these labels to reward logical consistency and penalize contradictions. Persona coverage, on the other hand, assesses how much of the given persona information is reflected in the response. This is captured by the Persona Distance (P-score or P-dist) [13], which computes the minimum cosine distance between the response embedding and persona sentence embeddings using a pretrained encoder. This enables detection of implicit personalization beyond exact lexical overlap.
To assess dialogue quality beyond personalization, we include two coherence metrics. The Utterance Entailment Score (UE-Score) [23] uses an NLI model fine-tuned on SNLI to evaluate whether each utterance is logically entailed by its preceding turn. The Coherence score from UniEval (Coh-UniEval) provides a model-based judgment of whether the response flows naturally within the broader dialogue context.
Together, these metrics offer a comprehensive and complementary evaluation: C-score and P-score target alignment with user-specific persona traits, while coherence metrics assess contextual flow and logical progression. A response can be coherent yet generic, fluent but ungrounded in the user’s identity, highlighting the limits of coherence alone. Traditional metrics like BLEU and ROUGE, based on $n$-gram overlap with fixed references, are ill-suited for personalization tasks where valid outputs are diverse and often lexically divergent. Our approach instead leverages fine-tuned language models and semantic embedding methods to enable reference-free, context-sensitive evaluation grounded in the user’s persona.
@@
LLMS: To implement an automatic evaluation pipeline, we chose the instructed versions of LLMs, as these models are fine-tuned to better follow structured prompts and format-specific instructions (e.g., JSON or reasoning), which is essential for ensuring consistency and reliability in automated output parsing. This also allows us to automate content extraction from the generated responses, streamlining the post-processing and evaluation modules. In this regard, we selected Mistral 7B [^7] https://mistral.ai/news/announcing-mistral-7b/, Qwen2 7B [^8] https://github.com/QwenLM/Qwen2, Gemma 7B [^9] https://huggingface.co/google/gemma-7b and Llama3.1 8B [^10] https://llama.meta.com/ as the four open-source LLMs that were executed on a computing device equipped with an NVIDIA RTX 6000 Ada Generation GPU. For the closed-source models, we utilized the GPT series from OpenAI [^11] https://openai.com/research/gpt, including GPT-3.5 Turbo, GPT-4 Turbo, and GPT-4o Mini, accessed through OpenAI’s official API and Gemini 1.5 Pro [^12] https://cloud.google.com/vertex-ai/generative-ai/docs/models/gemini/1-5-pro from Google AI.
5 Results and Analysis
@@
Table 4: Diversity evaluation using Dist-1 and Dist-2 metrics across the datasets in Vanilla and CoT setups.
Model IT-ConvAI2 BST FoCus
@@
To what extent can LLMs generate natural responses that are linguistically fluent? The benchmarked LLMs generally demonstrate strong fluency across various datasets, regardless of context length. This is particularly evident in the high metric values in Table 3, which demonstrate the response’s naturalness and the models’ ability to effectively capture the meaning and structure of conversations.
@@
Both groups perform well when comparing open-source and closed-source LLMs, with closed-source models like those in the GPT-4 series slightly outperforming their open-source counterparts, particularly in more extended contexts such as those in the FoCus dataset. However, the performance gap is not substantial, showing that open-source models are still capable, especially in less complex dialogues. Among open-source LLMs, Qwen2 performs well in shorter contexts, Gemma shows strong results across all datasets under the Vanilla setup but drops notably with CoT prompting, and Llama3.1 excels in longer, more complex dialogues, especially on the FoCus dataset. Meanwhile, GPT-4 models consistently demonstrate robust fluency across all contexts. Although CoT generally enhances their fluency in shorter contexts, it can sometimes introduce unnecessary complexity, slightly reducing performance by making responses more rigid and formulaic. Despite this, GPT-4 models maintain strong overall fluency in both short and long dialogues.
@@
How proficient are LLMs at generating diverse responses? Benchmarked LLMs are generally good at generating diverse responses, particularly in shorter-context datasets like IT-ConvAI2 and BST (Table 4). This is supported by high Dist-1 and Dist-2 scores, which measure the diversity of unigrams and bigrams, indicating a wide variety of words and varied word combinations in the generated responses. When comparing the closed-source and open-source LLMs, closed-source models like those in the GPT-4 series tend to perform slightly better overall, especially in longer contexts. However, the difference between the two groups is not stark, indicating that open-source models are still quite capable, particularly in less complex dialogues. For open-source LLMs, performance tends to decline as the context length increases, as seen in the FoCus dataset, which involves longer and more complex dialogues. This suggests that while these models handle diversity well in shorter exchanges, they face challenges maintaining the same diversity level in longer contexts. The impact of CoT prompting is most pronounced in open-source LLMs. Gemma shows a consistent performance drop on short-context datasets under CoT, yet both Gemma and Llama3.1 benefit significantly from CoT prompting in long-context evaluations. In contrast, closed-source models like GPT-4 show minimal improvement with CoT and sometimes even degrade in shorter contexts, as CoT may overstructure responses, reducing natural diversity. Nevertheless, GPT-4 models consistently excel across all context lengths, demonstrating strong performance in both short and long dialogues.
@@
Table 5: Personalization evaluation using P-Dist Score and C Score across the dataset, in Vanilla and CoT setups.
Model IT-ConvAI2 BST FoCus
@@
P-Dist C Score P-Dist C Score P-Dist C Score P-Dist C Score P-Dist C Score P-Dist C Score
@@
Table 6: Coherence evaluation using UE Score and UniEval Coherence metric (CH-UNE) on the datasets in Vanilla and CoT setups.
Model IT-ConvAI2 BST FoCus
Vanilla CoT Vanilla CoT Vanilla CoT
UE CH-UNE UE CH-UNE UE CH-UNE UE CH-UNE UE CH-UNE UE CH-UNE
Mistral 7B 0.46 0.84 0.31 0.83 0.48 0.93 0.44 0.94 0.21 0.56 0.33 0.90
Qwen2 7B 0.42 0.90 0.38 0.87 0.39 0.97 0.52 0.97 0.17 0.49 0.36 0.73
Gemma 7B 0.43 0.84 0.25 0.73 0.30 0.97 0.35 0.91 0.22 0.57 0.26 0.85
Llama3.1 8B 0.23 0.83 0.22 0.80 0.33 0.92 0.37 0.94 0.17 0.55 0.33 0.95
Gemini 1.5 Pro 0.45 0.87 0.33 0.90 0.46 0.93 0.44 0.97 0.26 0.75 0.32 0.83
GPT3.5 Turbo 0.49 0.90 0.38 0.87 0.51 0.97 0.57 0.96 0.37 0.91 0.45 0.93
GPT4o Mini 0.40 0.96 0.38 0.96 0.38 0.99 0.39 0.99 0.35 0.91 0.37 1.0
GPT4 Turbo 0.43 0.94 0.42 0.91 0.56 0.98 0.50 0.98 0.39 0.96 0.45 0.99
@@
How effectively can LLMs generate personalized responses that reflect the provided persona? The results of P-Dist and C scores, which are shown in Table 5, two well-established NLI-based metrics, reveal that LLMs struggle significantly with personalization. There is no clear benefit in using closed-source LLMs over open-source alternatives, underscoring the inherent difficulty of the personalization task for these models. Even in the best scenarios, the P-Dist score, where 1 represents the ideal outcome, demonstrates only moderate performance. The C score, which also penalizes contradictory entailments with negative values, exposes these limitations even more starkly, particularly in short-length contexts where less content is available for personalization. On short-context datasets, no clear performance trend emerges among open-source LLMs. However, on the FoCus dataset, Llama3.1 shows a slight superiority in the personalization task. The same overall scenario applies for closed-source LLMs, while GPT-4 Turbo emerges as the least effective performer. CoT negatively impacts short-length conversations but yields a notable improvement in long-length conversations, as seen with the FoCus dataset. This suggests that the added complexity of CoT is only advantageous when there is enough content to work with. This need for CoT reasoning aligns with findings from [29], which highlight that current language models do not reliably utilize information in long input contexts.
To what extent do LLMs generate coherent responses aligned with the dialogue context and persona? Based on the UE score and Coh-UniEval presented in Table 6, LLMs generally struggle with generating coherent responses, especially when working with long-context datasets like FoCus. Even in shorter-context datasets, as seen in the IT-ConvAI2 results, LLMs show poor performance in maintaining coherence. While the closed-source LLMs perform better at this task, even powerful models like GPT-4-Turbo fall short of achieving good results. Among them, GPT4o Mini demonstrates marginally lower performance across all the datasets with varying context sizes. While CoT reasoning offers a modest performance improvement, GPT series LLMs, as fine-tuned LLMs, remain limited in their ability to handle the coherence task effectively. Open-source LLMs exhibit closely aligned performance, with no clear outlier, suggesting that they collectively struggle to maintain coherence, especially in complex, long-context scenarios.
@@
Figure 2: Performance analysis of (a) open-source LLMs in vanilla setting, (b) closed-source LLMs in vanilla setting, (c) open-source LLMs in CoT setting, (d) closed-source LLMs in CoT setting on the FoCus dataset.
@@
Figure 3: Analysis of response time, failure ratio, and engagingness of LLMs on the FoCus dataset: (a) Vanilla and (b) CoT setups.
Table 7: LLM rankings on the FoCus dataset (Vanilla vs. CoT setups). Reported statistical tests indicate significance at $p<0.05$.
Vanilla CoT
Rank LLM Score Rank LLM Score
1 GPT-3.5 Turbo 50.69 1 Llama-3.1 8B 58.40
2 GPT-4 Turbo 50.13 2 GPT-4.0 Mini 57.43
3 GPT-4.0 Mini 49.83 3 GPT-3.5 Turbo 54.68
4 Gemini 1.5 Pro 40.28 4 GPT-4 Turbo 53.71
5 Gemma 7B 29.74 5 Mistral 7B 53.38
6 Mistral 7B 25.31 6 Gemma 7B 51.13
7 Llama-3.1 8B 24.30 7 Gemini 1.5 Pro 45.56
8 Qwen2 7B 21.29 8 Qwen2 7B 43.43
6 Discussion
To further analyse the LLM performance on longer text, we focused on the FoCus dataset. Our previous results show that its long-context nature makes LLM performance differences more apparent.
Fig. 2 provides an overall view of LLM groups across different prompting configurations. For visualization purposes, metrics with negative values were linearly transformed by adding a minimal constant offset to ensure positive values while preserving relative model comparisons. BertScore-F1, Dist-2, UE-Score, P-Dist, UniEval (Groundedness), and C Score are considered for overall fluency, diversity, coherence, persona coverage, persona consistency, and Groundedness analysis. BertScore-F1, Dist-2, UE-Score, P-Dist, UniEval (Groundedness), and C Score are considered for overall fluency, diversity, coherence, persona coverage, persona consistency, and Groundedness analysis. Response personalization remains challenging, even for fine-tuned closed-source LLMs, as evidenced by the results in both persona consistency and persona coverage. Groundedness, which evaluates the alignment of a response with the given conversational context, and Coherence, which ensures consistency with both the persona and the target response, further emphasize this challenge. While open-source LLMs generally cannot compete with closed-source models in terms of fluency and diversity, their performance in personalization, groundedness and coherence is considerably enhanced with the use of CoT. In fact, an open-source model like Llama3.1 can even surpass fine-tuned closed-source LLMs in these areas. Meanwhile, the performance of closed-source LLMs do not show any tangible improvement.
Fig. 3 illustrates LLMs’ comparative performance across response time, failure ratio, and engagingness, factors offering practical insights for selecting LLMs in industrial applications. The failure ratio is defined as the percentage of outputs (out of 1000) that failed JSON parsing, indicating how reliably a model follows structured output instructions. While CoT increases response time (measured as model generation time excluding any pre-/post-processing), it enhances instructability across both open- and closed-source LLMs. Although we observed a difference in how models handle token length limits, we did not conduct further quantitative analysis of this factor, and have included it as a limitation. Regarding engagingness, while an overall gap exists between model groups, CoT reasoning notably improves the perceived engagingness of open-source LLM outputs.
In Table 7, we present the ranking of benchmarked LLMs on the FoCus dataset under both the Vanilla and CoT setups. The reported scores are averaged over the evaluation metrics, Fluency, Diversity, Coherence, Persona Coverage, Groundedness, Persona Consistency, and Instructability. As shown, closed-source models such as GPT-3.5 Turbo and GPT-4 Turbo dominate in the Vanilla setup, while Llama-3.1 8B achieves the top performance under the CoT setup. [^13] https://github.com/good-researcher/PersoBench/blob/main/statistical_significance.ipynb